# Amorphous Boron Nitride as an Ultrathin Copper Diffusion Barrier for Advanced Interconnects

### v3

@@ 5 Experimental Investigation of -BN Diffusion Barriers
After we showed the potential of -BN films using theoretical models, we conducted an experimental investigation on the diffusion barrier performance of -BN films with the same thicknesses we performed our simulations. We fabricated Cu/-BN/Si structures and annealed them at 400 - 600 ℃ for 30 minutes. After annealing at 400 ℃, a Si substrate coated with a 3-nm thick -BN film did not form any copper silicide (Fig. 9-a), while a Si substrate not covered with -BN formed a substantial amount of copper silicide (Fig. 9-e). After annealing 3-nm thick -BN/Si at 500 ℃, very small copper silicide nucleations were observed on the Si surface (Fig. 9-b), and the formation of copper silicide crystals was confirmed after annealing at 600°C (Fig. 9-d). Fig. 9-c provides a brief summary of Cu diffusion barrier properties of 3-nm, 5-nm, and 7-nm thick -BN films after the 400 - 600 ℃ annealing process. The observation of crystallization by SEM is fine but not sensitive enough to determine Cu diffusion in the sub-silicide formation regime - that is - you could have Cu in the Si but SEM resolution is insufficient to detect or within the solubility limit of Cu in Si (Fig. 13 in C). The characteristics such as density or existence of defects of experimentally synthesized -BN films differ from the simulated -BN films and do not exhibit the same barrier performance as the simulations. However, there is a tendency for the barrier property to improve as the thickness of the -BN film increases.
@@ 6 Conclusions
We have investigated the diffusion barrier properties of ultrathin amorphous boron nitride (-BN) films against copper penetration by molecular dynamics and supported the simulations with experimental data. We focused on -BN films with thicknesses in the few nanometres range, and have analysed the morphologies, mechanical, diffusion barrier and dielectric properties of various -BN samples generated by molecular dynamics with machine learning trained interatomic potentials. The theoretical results agree with experimental barrier performance against Cu diffusion measured on PECVD-deposited 3-, 5- and 7-nm thick -BN films. Overall, the results confirm that amorphous boron-nitride compounds can be used as a dielectric barrier against Cu diffusion as a way to reduce the time delay in interconnects architectures. Further computational studies of -BN with larger thickness are however needed to explore (and find ways to mitigate) the effect of crystallisation observed in experiments, as it could could jeopardize reaching the upper performance and integrability of this innovative advanced material into mainstream interconnects technologies. Improved modelling of dielectric properties, beyond the qualitative tight-binding model used here, are likewise needed to better assess the properties of the films for such applications. In addition to a better electronic structure description, a more precise modelling of the dielectric function in the MHz to GHz range (including vibrational and intraband effects) would be of importance for comparison with experiments and to analyse the behaviour of -BN films at device operation frequencies.
@@ Acknowledgements
This project has been supported by Samsung Advanced Institute of Technology and is conducted under the REDI Program, a project that has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 101034328. This paper reflects only the author’s view and the Research Executive Agency is not responsible for any use that may be made of the information it contains. This work was supported by Institute for Basic Science (IBS-R036-D1), Republic of Korea. ICN2 acknowledges the Grant PCI2021-122092-2A funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR”. H.-J.S. acknowledge support from the National Research Foundation of Korea (NRF) grant funded by the Korea government Ministry of Science and ICT (RS-2024-00352458). Simulations were performed at the Texas Advanced Computing Center (TACC) at The University of Texas at Austin and the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DE-AC02-06CH11357.
@@ A.1 Generation of the Training Data
The GAP model used in this study was developed using two sets of samples, “training” and “validation” sets which include positions of atoms, atomic positions, energies, forces, and stresses calculated using Car Parrinello Molecular Dynamics and density functional theory (DFT) with Quantum Espresso package [41, 42, 43]. A multi-step procedure was employed to train the GAP model: 1) large number of amorphous structures were generated using Car-Parrinello MD, 2) the initial training set is generated using DFT calculations, 3) a GAP model is trained over this set; 4) a new set of validation set is generated using the previously generated GAP model; 5) the force and energy estimation of GAP model is evaluated using DFT calculations. These iterative training-validating steps are repeated until sufficiently low RMSE values are reached.
The training and validation sets together contain approximately 4500 structures. Here, the training set was constituted of atomistic structures from different phases of boron nitride, Cu-diffused -BN samples, -BN/Cu interfaces, and various Cu structures. The densities of all amorphous random samples (-B, -BN, -BxNyCuz, where x, y and z are random values between 0 and 1, and satisfy the condition ) are set to be between 1.5 - 5.0 . All amorphous random structures have cubic cells, with the size of the cell determined randomly. Crystalline BN phases (wurtzite, cubic, hexagonal) and Cu phases were (hcp, fcc, bcc, tetragonal) were taken from the Materials Project database[44]. We also introduced some defects and strain, as well as the melted liquid samples to ensure comprehensive coverage of potential configurations. Once the dataset was generated we randomly split it into training and validation sets which contain 2503 and 2000 structures, respectively. Periodic boundary conditions are applied to the cell for all structures. Ultrasoft pseudopotentials (USPP) with LDA as an exchange-correlation functional have been used for Car-Parrinello MD calculations. The kinetic energy cutoff for wavefunctions and energy are set to 150 Ry and 20 Ry. DFT calculations have been carried out using PAW pseudopotentials with PBE exchange-correlation functionals. Energy cutoffs of 75 Ry and 600 Ry have been employed for the wavefunction and electron density, respectively. To account for the van der Waals (vdW) interactions between BN/Cu interfaces, the non-local vdw-df-obk8 functionals[45] were used. A Gaussian smearing of 0.1 eV width is applied to electronic levels. Car-Parrinello MD simulations were performed using -point sampling, while DFT calculations were performed using Monkhorst-Pack grids of 8x8x8 for 3D samples and 8x8x1 for monolayer samples.
@@
The GAP model is trained using the QUIP software [46, 20]. We employed three descriptors: two-body(2b), three-body(3b), and smooth overlap of atomic positions (SOAP). The 2b descriptor is used to describe the pairwise interactions between atoms. The 3b and SOAP descriptors represent the contribution of the atomic triplets and the local environment around an atom. The cut-offs and hyperparameters for each descriptor have been presented in Table 1. The GAP models have been proven to be useful for several materials [18, 19, 22, 47]. Sparsification using the CUR method [48] was selected for the SOAP kernel, while sparsification on a uniform grid was used for two-body and three-body kernels. The quality of the trained GAP model was later assessed by comparing the obtained energies per atom and forces of samples in the training and validation sets with the trained GAP model and DFT calculations. The root-mean-square errors for energy and forces on training data are 0.0029 eV/atom and 0.296 eV/ and on validation data are 0.0031 eV/atom and 0.306 eV/, as shown in Fig. 10. The obtained RMSE values are similar to the values in the literature for similar amorphous structures [49, 22, 50]. Given that amorphous materials and complex systems, like the one we presented this study, require a highly diverse dataset, the RMSE values are considered as acceptable[50].
@@ Appendix B Additional Morphological Information
Appendix C SEM Images of of -BN Diffusion Barriers

### v4

@@ 5 Experimental Investigation of -BN Diffusion Barriers
After we showed the potential of -BN films using theoretical models, we conducted an experimental investigation on the diffusion barrier performance of -BN films with the same thicknesses we performed our simulations. We fabricated Cu/-BN/Si structures and annealed them at 400 - 600 ℃ for 30 minutes. After annealing at 400 ℃, a Si substrate coated with a 3-nm thick -BN film did not form any copper silicide (Fig. 9-a), while a Si substrate not covered with -BN formed a substantial amount of copper silicide (Fig. 9-e). After annealing 3-nm thick -BN/Si at 500 ℃, very small copper silicide nucleations were observed on the Si surface (Fig. 9-b), and the formation of copper silicide crystals was confirmed after annealing at 600°C (Fig. 9-d). Fig. 9-c provides a brief summary of Cu diffusion barrier properties of 3-nm, 5-nm, and 7-nm thick -BN films after the 400 - 600 ℃ annealing process. The observation of crystallisation by SEM is fine but not sensitive enough to determine Cu diffusion in the sub-silicide formation regime - that is - you could have Cu in the Si but SEM resolution is insufficient to detect or within the solubility limit of Cu in Si (Fig. 16 in D). The characteristics such as density or existence of defects of experimentally synthesized -BN films differ from the simulated -BN films and do not exhibit the same barrier performance as the simulations. However, there is a tendency for the barrier property to improve as the thickness of the -BN film increases.
@@ 6 Conclusions
We have investigated the diffusion barrier properties of ultrathin -BN films against copper penetration by molecular dynamics and complemented the simulations with experimental data. Simulations show that Cu diffusivity is strongly suppressed in dense, slowly quenched films with low concentrations of B–B and N–N bonds, reduced content, and high atomic packing. These structures remain effective up to high temperatures, while rapidly quenched, defect-rich networks permit Cu diffusion at lower temperatures due to increased homonuclear bonding and reduced structural cohesion. Increasing the film thickness enhances performance by enabling structural relaxation and lowering defect concentrations. These trends are consistent with experimental measurements on PECVD-deposited 3-, 5-, and 7-nm films, which show barrier failure temperatures that increase with thickness. A central experimental challenge, however, is to preserve a controlled amorphous structure in BN films as their thickness increases, since B-N rings intrinsically tend to align and nucleate turbostratic or nanocrystalline h-BN domains [17]. At small thicknesses, limited adatom mobility and kinetic constraints favour the stabilisation of a disordered network. However, as growth proceeds, enhanced surface and bulk diffusion, stress relaxation, cumulative plasma energy input, and local deviations from the ideal B:N stoichiometry progressively promote the emergence of ordered regions. This structural evolution not only undermines the targeted amorphousness but also modifies critical functional properties – including dielectric constant, density, and diffusion-barrier performance – thus making crystallisation control a key bottleneck for scaling -BN films in interconnect and dielectric applications. To advance in this direction, it would be valuable to complement experiments with computational insights, for instance on the dependence of the dielectric constant on growth rate [15], or on the roles of growth temperature and oxygen incorporation [14]. Further computational studies of -BN with larger thickness are also needed to explore (and find ways to mitigate) the effect of crystallisation observed in experiments, as it could jeopardise reaching the upper performance and integrability of this innovative advanced material into mainstream interconnect technologies. Improved modelling of dielectric properties, beyond the simplified tight-binding model used here, are likewise needed to better assess the properties of the films for such applications. In addition to a better electronic structure description, a more precise modelling of the dielectric function in the MHz to GHz range (including vibrational and intraband effects) would be of importance for comparison with experiments and to analyse the behaviour of -BN films at device operation frequencies. Overall, the results confirm that amorphous boron nitride compounds are effective dielectric barriers against Cu diffusion, offering a promising strategy to reduce time delays in scaled interconnect architectures.
@@ Acknowledgements
This project has been supported by Samsung Advanced Institute of Technology and is conducted under the REDI Program, a project that has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 101034328. This paper reflects only the author’s view and the Research Executive Agency is not responsible for any use that may be made of the information it contains. This work was supported by Institute for Basic Science (IBS-R036-D1), Republic of Korea. ICN2 acknowledges the Grant PCI2021-122092-2A funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR”. S. R. is also supported by MICIN with European funds NextGenerationEU (PRTRC17.I1) funded by Generalitat de Catalunya and by 2021 SGR 00997. H.-J.S. acknowledge support from the National Research Foundation of Korea (NRF) grant funded by the Korea government Ministry of Science and ICT (RS-2024-00352458). Simulations were performed at the Texas Advanced Computing Center (TACC) at The University of Texas at Austin and the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DE-AC02-06CH11357.
@@ A.1 Generation of the Training Data
The GAP model used in this study was developed using two datasets, referred to as the “training” and “validation” sets, which include atomic positions, energies, forces, and stresses calculated using Car-Parrinello molecular dynamics (CPMD) and density functional theory (DFT) with the Quantum ESPRESSO package [46, 47, 48]. The training process followed an iterative scheme: (1) a large number of initial structures were generated by calculating the energies, forces, and stresses of molecular compounds, crystalline phases, and amorphous structures containing B, N, and Cu atoms; (2) an initial GAP model was trained on this dataset; (3) a new set of validation structures was generated using the previously trained GAP model; (4) the energy and force predictions of the GAP model were evaluated using DFT calculations. These training–validation cycles were repeated until the lowest RMSE values were achieved. Additionally, the accuracy of the final model was further assessed by comparing structural and energetic properties, as described below.
The “training” set used to generate and validate the final GAP model contains approximately 2500 structures, while the final validation set comprises 2000 structures. Together, the dataset includes around 200,000 atomic environments. A detailed breakdown by system type, number of atoms per configuration, and number of configurations is provided in Table A.1. The dataset spans a wide variety of configurations to ensure comprehensive coverage of relevant atomic environments. Initially, we incorporated several molecular compounds of B and N, as well as Cu systems, available in the Materials Project database [49]. Crystalline forms of BN and Cu were also included using the same source. These crystalline samples were expanded into larger supercells, and additional configurational diversity was introduced through the application of random defects and strain. Crystalline BN/Cu heterostructures were also generated to model interfacial environments. To cover disordered regimes, we generated a large number of liquid and quenched structures. These include -B, -Cu, -BN, -BN/Cu heterostructures, and randomly generated -BxNyCuz configurations, where , , and are random values constrained by . All amorphous and liquid samples contain 64 atoms per unit cell, with cubic cell geometries. Cell parameters were randomly assigned to ensure that the resulting densities span the range 1.5–5.0 g/cm3. Periodic boundary conditions were applied to all systems. For layered structures and heterostructures, a 10 vacuum spacing was added along the -direction; for isolated atoms and molecular compounds, vacuum was added in all three spatial directions.
| System | Num. of Atoms | Num. of Config. |
| Isolated elements (B, N, Cu) | 1 | 3 |
| Small molecules (B2, N2) | 2–16 | 100 |
| Crystalline phases (e.g.,h-BN, c-BN, fcc-Cu) | 4–64 | 1900 |
| Amorphous structures (e.g., -BN, -BxNyCuz) | 64 | 2500 |
Breakdown of the training and validation datasets by system type, number of atoms per configuration, and total number of configurations used for GAP model development.
Car-Parrinello MD simulations for melt and quenched samples were performed using ultrasoft pseudopotentials (USPP) with the LDA exchange-correlation functional, employing kinetic energy cutoffs of 20 Ry for wavefunctions and 150 Ry for the charge density. DFT calculations for dataset labelling were performed using PAW pseudopotentials with the PBE exchange-correlation functional, with plane-wave cutoffs of 75 Ry for wavefunctions and 600 Ry for the charge density. To describe van der Waals interactions at BN/Cu interfaces, the non-local vdw-df-obk8 functional was employed [50]. When structures were relaxed, ionic steps were continued until all atomic forces were below 0.001 atomic units (a.u.). A Gaussian smearing of 0.1 eV was applied to the electronic levels during DFT calculations. -point sampling was used for CPMD simulations, while Monkhorst-Pack grids (888 for three-dimensional systems and 881 for monolayers) were employed for DFT calculations.
@@
The GAP model was trained using the QUIP software [51, 20]. We employed three descriptors: two-body(2b), three-body(3b), and smooth overlap of atomic positions (SOAP). The 2b descriptor is used to describe the pairwise interactions between atoms. The 3b and SOAP descriptors represent the contribution of the atomic triplets and the local environment around an atom. The cut-offs and hyperparameters for each descriptor have been presented in Table 1. The GAP models have been proven to be useful for several materials [18, 19, 24, 52]. Sparsification using the CUR method [53] was selected for the SOAP kernel, while sparsification on a uniform grid was used for two-body and three-body kernels. The quality of the trained GAP model was later assessed by comparing the obtained energies per atom and forces of samples in the training and validation sets between the GAP model predictions and DFT calculations. The root-mean-square errors for energy and forces on training data are 0.0029 eV/atom and 0.296 eV/ and on validation data are 0.0031 eV/atom and 0.306 eV/, as shown in Fig. 10. The obtained RMSE values are similar to the values in the literature for similar amorphous structures [54, 24, 55]. Given that amorphous materials and complex systems, like the one we presented this study, require a highly diverse dataset, the RMSE values are considered as acceptable[55].
@@ Appendix B Additional Morphological Information
Appendix C Diffusivity vs Film Mass Density at High Temperatures
Appendix D SEM Images of -BN Diffusion Barriers